I do not understand can you help?
Simplify the problem and write answers in exponential form.
3^4 + 2 * 3^4

Answer by stanbon(48504)

I do not understand can you help?
Simplify the problem and write answers in exponential form.
3^4 + 2 * 3^4
---
(3^4) + 2(3^4)
---
= (3^4)(1 + 2)
---
= (3^4)*3
---
= 3^5
= 243
==========
Cheers,
Stan H.